# GOAL: A Generalist Combinatorial Optimization Agent Learning

### v2

@@ Appendix C Impact of codebook on fine-tuning
Appendix D Generalization to new instance distributions
@@
Train. distr. Generalization
ATSP100 ATSP200 ATSP500 ATSP1000
MatNet greedy 1.57% 1s 97.21% 12s - -
BQ-NCO greedy 1.63% 1s 1.90% 9s 4.29% 26s 8.09% 1m
GOAL multi-task greedy 0.50% 5s 1.84% 53s 1.77% 3m 2.37% 28m
CVRP 100 CVRP200 CVRP500 CVRP1000
AM greedy 7.11% 1s 10.16% 1s 17.2% 1s 72.95% 2.1s
MDAM greedy 4.84% 1s 7.54% 6s 12.94% 24s 32.25% 1m
POMO no aug 1.21% 1s 5.94% 6s 31.36% 32s 284.58% 1m
Sym-NCO greedy 3.33% 1s 8.42% 3s 37.58% 54s 481.53% 9m
BQ-NCO greedy 2.79% 1s 2.81% 9s 3.64% 55s 5.88% 2m
GOAL multi-task greedy 3.06% 5s 3.35% 30s 3.96% 2m 7.63% 3m
Table 2: Generalization performance on larger instances. Lower is better.
Appendix E Description of the considered CO problems
@@
E.1 Capacitated Vehicle Routing Problem and Split-Delivery Capacitated Vehicle Routing Problem
@@
Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10]. Following [kool_attention_2019], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
@@
E.2 Capacitated Vehicle Routing Problem with Time Windows
@@
E.3 Job Shop Scheduling Problem
The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem. We are given $N$ jobs and $M$ machines. Each job consists of a sequence of operations (they can called tasks) that need to be processed in a specific order (implying precedence constraints). Each operation has a given processing time. Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption). A solution is a schedule, defined by the start time of each operation such that a specific criterion, such as makespan, total completion time, or total tardiness is minimized.
@@ E.3 Job Shop Scheduling Problem
Node features: Job execution time, job availability time, machine availability time
Edge features: Precedence constraint, task-machine dependency
@@
Action: Select a task
@@ E.4 Knapsack Problem
E.5 Maximal Independent Set Problem
@@
E.6 Maximum Coverage Location Problem
@@
E.7 Minimum Vertex Cover Problem
@@
E.8 Open Shop Scheduling Problem
@@
E.9 Open Vehicle Routing Problem
Open Vehicle Routing Problem (OCVRP) is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers. In contrast with other CVRP variants, vehicles do not need to return to the depot at the end of their service. This makes OCVRP similar to the classical CVRP problem, with a different objective function. In our experiments, we handle the Euclidean version of the OCVRP and fine-tune GOAL to solve it.
Data generation.
Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10]. As in CVRP, the vehicle capacity is set to 50 for a problem size of 100.
Instance features: Remaining vehicle capacity
Node features: Origin/destination tokens, demand
Edge features: Distance between nodes
Oracle: LKH
Objective: Minimize traveled distance
Action: Select a node, with direct edge or via-depot
remaining capacity in descending order
E.10 Orienteering Problem
The Orienteering Problem (OP) is a combinatorial optimization problem that involves finding the optimal route to maximize the collected prizes while visiting a set of locations and satisfying a maximal distance constraint.
Data generation.
We generated data following previous work [kool_attention_2019] and  [xin_multi-decoder_2021]: node locations are uniformly sampled from the unit square and prizes are proportional to the distance to the depot. We followed the same procedure line in  [kool_attention_2019], this variant of the problem is called $OP_{distance}$. The maximal tour length constraint is set to 4. We use Euclidean OP for training GOAL.
Instance features: Remaining tour length
@@ E.10 Orienteering Problem
E.11 Price Collection Traveling Salesman Problem
@@
E.12 Sequential Ordering Problem
@@
E.13 Traveling Repairman Problem
@@
E.14 Traveling Salesman Problem
@@
For the Euclidean case, we generate datasets as described in [kool_attention_2019]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed. For non-Euclidean problems (ATSP) we generate data as described in [kwon_matrix_2022], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
@@
E.15 Unrelated Machine Scheduling Problem

### v3

@@ Appendix C Impact of codebook on fine-tuning
Appendix D Multi-type mixed attention layers for UMSP and JSSP
As an illustration of our proposed multi-type transformer architecture, we present in Figure 7 the inputs to the mixed-attention blocks for the Unrelated Machine Scheduling (UMSP) and the Job Shop Scheduling Problems (JSSP). Note that in the case of UMSP, since there is no edge information between jobs and between machines, the self-attention blocks (left-most and right-most attention blocks) degenerate to vanilla attention. The cross-attention blocks, on the other hand, are mixed, and they are fed the same edge information, holding the job execution times on each machine (transposed from each other). In JSSP, the situation is similar, but ops have two kinds of relations between themselves: precedence, and sharing the same job. Hence the self-attention block on ops is mixed.
(a) the Unrelated Machine Scheduling Problem (UMSP)
(b) the Job-shop Scheduling Problem (JSSP)
Figure 7: Illustration of the multi-type model on scheduling tasks in GOAL. The "MA" blocks use vanilla multi-head attention while the "MMA" use our multi-head mixed-attention.
Appendix E Generalization to new instance distributions
@@
Train. distrib. Generalization
100 200 500 1000
ATSP LKH 0.00% 29s 0.00% 48s 0.00% 2m 0.00% 4m
MatNet 1.57% 1s 97.21% 12s - -
BQ-NCO 1.63% 1s 1.90% 9s 4.29% 26s 8.09% 1m
GOAL 0.45% 5s 1.54% 30s 1.62% 3m 1.96% 5m
CVRP LKH 0.00% 12m 0.00% 2h 0.00% 5h 0.00% 7h
AM 7.11% 1s 10.16% 1s 17.2% 1s 72.95% 2s
MDAM 4.84% 1s 7.54% 6s 12.94% 24s 32.25% 1m
POMO 1.21% 1s 5.94% 6s 31.36% 32s 284.58% 1m
Sym-NCO 3.33% 1s 8.42% 3s 37.58% 54s 481.53% 9m
BQ-NCO 2.79% 1s 2.81% 9s 3.64% 55s 5.88% 2m
GOAL 3.16% 5s 3.42% 30s 3.90% 2m 7.37% 5m
CVRPTW HGS (1s) $\dagger$ 108/128 2m 14/128 2m 0/128 2m 0/128 2m
HGS (1m) 0.00% 10m 0.00% 26m 0.00% 2h 0.00% 2h
GOAL 3.82% 5s 7.71% 30s 9.49% 2m 6.75% 5m
OP EA4OP 0.00% 1m 0.00% 3m 0.00% 11m 0.00% 45m
AM 5.35% 1s 11.05% 1s 22.29% 1s 30.73% 1s
MDAM 2.88% 16s 5.55% 1s 18.31% 4s 30.05% 15s
SymNCO 2.03% 2s 6.60% 3s 19.81% 24s 31.79% 1m
GOAL 0.43% 3s 1.60% 10s 7.45% 17s 15.62% 30s
KP ORTools 0.00% $<1$s 0.00% $<1$s 0.00% $<1$s 0.00% $<1$s
POMO 0.19% 1s 0.50% 1s 6.41% 31s 5.34% 1m
BQ-NCO 0.10% 2s 0.14% 6s 0.74% 7s 0.92% 13s
GOAL 0.12% 3s 1.63% 6s 2.40% 10s 2.59% 16s
10x10 10x15 15x15 20x15
JSSP ORTools 0.00% 47s 0.00% 2m 0.00% 3h 0.00% 80h
COMPASS $\diamond$ 4.70% 3h - 8.00% 5h 10.40% 8h
GOAL 4.13% 15s 11.02% 43s 10.74% 45s 22.53% 4m
Table 2: Generalization performance on larger instances. Lower is better. All NCO methods perform greedy decoding, except for methods marked with $\diamond$; they do not provide results for greedy decoding. Solvers marked with $\dagger$ are unable to solve all instances within the given time budget; we report the number of instances successfully solved.
Appendix F Description of the considered CO problems
@@
F.1 Capacitated Vehicle Routing Problem and Split-Delivery Capacitated Vehicle Routing Problem
@@
Node locations are randomly sampled from the unit square, while node demands are randomly generated integers from the interval [1, 10]. Following [kool_attention_2018], the vehicle capacity depends on the problem size: capacity of 50 for a problem size of 100, 80 for 200, 100 for 500, and capacity of 250 for problems with 1000 nodes.
@@
F.2 Capacitated Vehicle Routing Problem with Time Windows
@@
F.3 Job Shop Scheduling Problem
The Job Shop Scheduling Problem (JSSP) is a classic scheduling problem. We are given $N$ jobs and $M$ machines. Each job consists of a sequence of operations that need to be processed in a specific order (implying precedence constraints). Each operation has a given processing time. Each machine can only process one operation at a time and when an operation is started it cannot be interrupted (no preemption). A solution is a schedule, defined by the start time of each operation such that a specific criterion, such as makespan, total completion time, or total tardiness is minimized.
@@ F.3 Job Shop Scheduling Problem
Node features: Operation execution time, job availability time, machine availability time
Edge features: Precedence constraint, operation-machine dependency
@@
Action: Select an operation
@@ F.4 Knapsack Problem
F.5 Maximal Independent Set Problem
@@
F.6 Maximum Coverage Location Problem
@@
F.7 Minimum Vertex Cover Problem
@@
F.8 Open Shop Scheduling Problem
@@
F.9 Open Vehicle Routing Problem
Open Vehicle Routing Problem (OCVRP) is a variant of CVRP where an unbounded fleet of vehicles is available at the depot, each serving a set of customers. In contrast with other CVRP variants, vehicles do not need to return to the depot at the end of their service. This makes OCVRP similar to the classical CVRP, as both problems involve a set of locations with demand and vehicle capacity constraints. However, the key difference lies in the route structure: in OCVRP, vehicles do not need to return to the depot after completing deliveries, which affects the solution space and cost minimization strategy. OCVRP has common applications in logistics, where delivery trucks do not need to return to the depot after completing deliveries. Data generation and problem features are the same as for the CVRP. In our experiments, we handle the Euclidean version of the OCVRP and fine-tune GOAL to solve it.
F.10 Orienteering Problem
The Orienteering Problem (OP) is a combinatorial optimization problem that combines aspects of the Traveling Salesman Problem (TSP) and the Knapsack Problem. It is commonly used in routing and scheduling applications where the goal is to maximize collected rewards while respecting a constraint on total travel distance or time. The OP is defined by a set of locations (nodes), each associated with a reward; a starting location and an ending location (usually the same node); a travel cost matrix between locations; and a maximum allowable travel budget. The objective is to visit a subset of locations that maximizes the total collected reward while ensuring the total travel cost does not exceed the budget. We generated data following previous work [kool_attention_2018]: node locations are uniformly sampled from the unit square and prizes are proportional to the distance to the depot (in  [kool_attention_2018] this variant of the problem is called $OP_{distance}$). The maximal distance constraint is set to 4. We use Euclidean OP for training GOAL.
Instance features: Remaining distance
@@ F.10 Orienteering Problem
F.11 Price Collection Traveling Salesman Problem
@@
F.12 Sequential Ordering Problem
@@
F.13 Traveling Repairman Problem
@@
F.14 Traveling Salesman Problem
@@
For the Euclidean case, we generate datasets as described in [kool_attention_2018]: points are uniformly sampled from the unit square, and the corresponding distance matrix $A$ is computed. For non-Euclidean problems (ATSP) we generate data as described in [kwon_matrix_2021], where distance matrices are generated randomly, but post-processed to satisfy the triangular inequality.
@@
F.15 Unrelated Machine Scheduling Problem